OpenStudy (anonymous):

SIMPLIFY?? 3√45 - 8√20

OpenStudy (anonymous):

@jim_thompson5910 helpp?(:

OpenStudy (dbzfan836):

-15.6524758425

jimthompson5910 (jim_thompson5910):

how would you simplify \(\Large \sqrt{45}\)

OpenStudy (anonymous):

@jim_thompson5910 you can't right? but the square root of 45 is 6.70

jimthompson5910 (jim_thompson5910):

can you factor 45 into two numbers where one of them is a perfect square?

OpenStudy (anonymous):

@jim_thompson5910 yes 9 and 5, 9 would be the perfect square

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so \[\Large \sqrt{45} = \sqrt{9*5}\]

OpenStudy (anonymous):

@jim_thompson5910 okay..

jimthompson5910 (jim_thompson5910):

then you use the rule \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{45} = \sqrt{9*5}\] \[\Large \sqrt{45} = \sqrt{9}*\sqrt{5}\] \[\Large \sqrt{45} = ???\]

OpenStudy (anonymous):

@jim_thompson5910 i don't know :(

OpenStudy (anonymous):

3√45 - 8√20 = 3√9*5 - 8√4*5 = 9√5 - 16√5 = -7√5 I go this from Yahoo Answers, does it help? It helped me. :)

jimthompson5910 (jim_thompson5910):

Do you see how I got \[\Large \sqrt{45} = \sqrt{9*5}\] \[\Large \sqrt{45} = \sqrt{9}*\sqrt{5}\]

OpenStudy (anonymous):

@jim_thompson5910 yes I see you used the equation from above..

jimthompson5910 (jim_thompson5910):

what is the square root of 9

OpenStudy (anonymous):

@CrazyCountryGirl thanks but I need a little more in depth help. :)

OpenStudy (anonymous):

@jim_thompson5910 3 is the square root of 9

jimthompson5910 (jim_thompson5910):

so, \[\Large \sqrt{45} = \sqrt{9*5}\] \[\Large \sqrt{45} = \sqrt{9}*\sqrt{5}\] \[\Large \sqrt{45} = 3*\sqrt{5}\]

OpenStudy (anonymous):

OK. :) Sorry I couldn't help more.

OpenStudy (anonymous):

@jim_thompson5910 oh okay i see..

jimthompson5910 (jim_thompson5910):

so how would you simplify \[\Large \sqrt{20}\]

OpenStudy (anonymous):

it would be 5 and 4 where 4 is the perfect square. and it would be √20= √5*4

jimthompson5910 (jim_thompson5910):

so far so good

OpenStudy (anonymous):

√20 = √5 √4

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

@jim_thompson5910 so it would be √20 = 2*√5

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

summing up we have so far \[\Large \sqrt{45} = 3*\sqrt{5}\] \[\Large \sqrt{20} = 2*\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

those two facts allow us to go from \[\Large 3*\sqrt{45} - 8*\sqrt{20}\] to \[\Large 3*3*\sqrt{5} - 8*2*\sqrt{5}\]

OpenStudy (anonymous):

@jim_thompson5910 so it would be 9 * √ 5 - 16 * √ 5

jimthompson5910 (jim_thompson5910):

good, what's next?

OpenStudy (anonymous):

then you subtract 9-16 and it would be -7√5

OpenStudy (anonymous):

@jim_thompson5910 thank you so much. :) saved my life.. haha

jimthompson5910 (jim_thompson5910):

you're welcome